Thermal decomposition of a compound is first order. If 50% of a sample of the compound is decomposed in 120 minutes, how long will it take for 90% of the compound to decompose?

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

half-life=ln(2)/(decay constant), 120=ln(2)/(lambda), solve for lambda.

10=100*e^-(lambda*t), t is the answer

somy chaturvedi
12 years ago

for 1 st order    for half concentration  k=0.693/t

                                                           k=0.693/120

                                                             now applty Ist order kinetics    t = 2.303/t*log a/a-xand get ur answer